OpenStudy (anonymous):

a student scores 65 and 82 out of 100 in the two test different test. How much does the student need to score in the next test which also has a max score of 100, in order that the average of the three test is a least 75?

OpenStudy (anonymous):

Since the average should be at least 75, the inequality sign to be used is

OpenStudy (anonymous):

greater or equal to sign

OpenStudy (anonymous):

How do you get an average? You add all the scores, divide it by 3 One of the scores will be x

OpenStudy (texaschic101):

(65 + 82 + x )/3 >= 75 multiply the equation by 3 65 + 82 + x >= 225 147 + x >= 225 x >= 225 - 147 x >= 78 the student has to make at least 78

OpenStudy (anonymous):

(65+82+x)/3≥75 miltiply both side by 3 65+82+x≥225 147+x≥225 x=225-147 x=78